While building your first telescope, your nerd friend suggested a brilliant idea for you to try and replace the reflecting mirror of the telescope by a fisk of mercury rotating about a vertical axis with some angular velocity w that forms a spherical surface of radius R, in such a telescope, the the focal length of the telescope would be function of the rotating speed of the mercury disk, calculate how much angular velocity is needed for the telescope to have a focal length of 15 cm?



1
Expert's answer
2021-05-11T16:54:21-0400

Focal length of the telescope (f)=15cm

So, radius of the curved surface "(R)=2f = 30cm"

Angular velocity "(V)= R\omega"

Or we can write it as "(R)=\frac{V}{\omega}"

"\Rightarrow 2f= \frac{V}{\omega}"


"\Rightarrow 30=\frac{V}{\omega}"